# thanks: These authors contributed equallythanks: These authors contributed equallythanks: These authors contributed equallythanks: These authors contributed equally

### v1

@@ VII A Model of a Kohn ladder
We consider a model calculation in which the temperature dependent self-energy is a scaled version of $\Pi_{33}$ in Fig. 2(b). To simulate anticrossing, we assume that the self-energy is initially concentrated entirely on the highest mode, $\Pi_{0,88}=\alpha\Pi_{33}$, where $\Pi_{0}$ is the bare self-energy and $\alpha$ increases as $T$ decreases. Figure 6 presents illustrative results, which are based on taking the bare phonons from the one-band model solutions for $\sigma=0.00565$ Ry. The results illustrate the anticrossing phenomena, and are in general accord with the QE results (green dashed line in (b)). Figure 6(a) uses the clean-limit $\Pi_{33}$, showing that the anomalous nesting peak spoils the agreement with experiment, as expected, while the dirty limit result, Fig. 6(b), restores the good agreement. Fig. 6(b) also show that for $\alpha\leq 5$, most of the softening is in the LO modes 6 and 8, while the LA mode barely changes. For larger $\alpha$, most of the softening is in the LA mode. The latter regime resembles the QE data for $\sigma\leq 0.03$ Ry.
@@
Figure 6: (color online) Model calculations of the evolution of the dressed frequencies at a series of temperatures scaled by $\alpha$. For $\alpha=0$ (highest $T$): one-band solutions for $\omega_{q}$ calculated from $\sigma=0.03$ Ry (black dotted lines); adding off-diagonal $\Pi$, with $|\Pi_{ij}|^{2}=20$ (blue lines). Lower temperatures: $\alpha$ =5 (violet lines), 8 (red lines), 9 (brown lines). Other curves: green dashed line = band 3 from $\sigma=0.07$ Ry data; thick dark red line = band 8 for $\alpha=9$ but $\Pi_{ij}=0$.
@@
Notably, if we turn off the mode-coupling, the softening of mode 8 at $\alpha=9$ (thick dark red curve) is too weak for a CDW instability. This can be understood as follows. Considering two modes only, the anti-crossing results in the renormalized frequencies given by:
@@
where the $\Omega_{1,2}$ are the renormalized frequencies found from Eqs.  1 (without mode mixing) and $\Omega_{\pm}$ include mode mixing. Clearly, $\Omega^{2}_{-}$ can go soft completely at the critical mode mixing threshold $\Pi_{12}^{(c)}$ where
@@
which takes place before $\Omega_{1}^{2}$ goes soft (without mode-mode coupling). Off-diagonal self-energies thus generally enhance $T_{CDW}$. Mode-mode mixing can also induce a shift in $\boldsymbol{Q}_{CDW}$. These effects can be shown for NbSe2 by investigating the difference between $\Omega_{\pm}$ and $\Omega_{1,2}$. Note that while the diagonal self-energies of Eq. 1 are real, the off-diagonal terms can be complex. In the irreducible representation with dimension 3, the Hamiltonian determinant contains the term $-16\omega_{1}\omega_{2}\omega_{3}Re(\Pi_{12}\Pi_{23}\Pi_{31})$. That is, the phonon energies depend on the relative phases between the off-diagonal components of the self-energy matrix. In contrast, for irreducible representations of lower dimension, the off-diagonal term only enters as $|\Pi_{12}|^{2}$.
@@
Interestingly, the initial electron-phonon interaction at high temperatures is longitudinal, so that the main phonon softening predominantly affects longitudinal modes. However, due to anticrossing, the phonon that softens will appear to be from a transverse mode, unless, as in NbSe2, the softening is along a high-symmetry line or point where longitudinal and transverse modes are strictly orthogonal.
VIII Discussion
VIII.1 Giant Kohn anomalies
The approach discussed in this paper would be applicable for understanding giant Kohn anomalies segmented into ladders via mode anticrossings in many correlated $d-$ and flimit-$f-$electron materials. Not including effects of anticrossing cloud underestimate the strength of electron-phonon coupling. Even in graphene, DFT calculations fail to properly capture the Kohn anomaly[graphene].
@@ VIII.1 Giant Kohn anomalies
A study of spin-lattice relaxation argues that most phonons play no role in relaxation, which is dominated by phonons with rotational or intramolecular vibrational patterns.[SLR1] Moreover, even these latter phonons lie at high energies, and onlt the small tail of these phonons at low energies dominates the relaxation. Our study suggests that phonons involved in relaxation are all optical, and that the low energy tail is produced by strong electron-phonon interaction in the form of Kohn anomalies, which push some modes to lower energies (soft modes). A chiral Kohn anomaly has indeed been reported recently in a Weyl semimetal.[Mingda]
@@
Figure 7: (color online) K-point phonon displacements. (a) mode 6, showing three phases of displacement in one cycle; (b) mode 7, and (c) mode 2 showing a single phase; and (d) phonon dispersions with red dots showing $k$-points used, from top to bottom mode 6 (a), 7 (b), and 2(c). Yellow boxes highlight one triangle, while violet boxes highlight one hexagon.
VIII.2 Visualizing soft phonons
@@
By using the phonon visualizer software [phonon_visualizer], we find that the phonon modes in NbSe2 with significant electron-phonon couplings undergo displacements that are circularly or elliptically polarized, raising interesting question as to what happens when a circular phonon softens. For a linearly polarized phonon, the frozen phonon CDW will have the maximum distortion associated with the soft phonon. But circular phonons have both amplitude and phase as variables, and the phase variable could lock in uniformly, adopt spatially random fixed directions (as in order-disorder transitions[OD]), or remain free as in a time crystal[time1, time2]. There is also a striking similarity between the circular phonon displacements and the dynamic Jahn-Teller (dJT) phenomena, which have been suggested as a source for time crystals.
We consider two $q$-values: $\bf q=\overrightarrow{\Gamma K}$, which lies in the region where the longitudinal-transverse anticrossing occurs and $\bf q=2/3\overrightarrow{\Gamma M}$, where the soft phonon lies. These two $q$ points are special in that they are 1/3 of a combination of reciprocal lattice vectors. Remarkably, most of the $K$-point phonons are associated with recognizable dJT displacements. One of the earliest dJT systems involves triangular molecules, such as Na3, which can have a Berry phase that leads to an extra orbital angular momentum with half-integer quantization (a ‘time molecule’).[LongHi, Na3] Similar phenomena can be found on impurity sites in other crystals.
@@
A layer of NbSe2 can be viewed as a lattice of centered triangular molecules, either triangles of Se2 molecules with a central Nb or of Nb molecules with a central Se2. For several $K$-point circular phonons one cycle of rotation amounts to the central atom displacing sequentially toward each of the triangular atoms, leading to a time-averaged symmetrical pattern (yellow squares in Fig. 7), as in Refs. LongHi, Na3. However, a lattice of dJT molecules has extra degrees of freedom, which also have a dJT interpretation. On each of the three steps in a cycle, a pattern forms on one set of atoms, then dissolves and reforms on a second and then a third set before it repeats in the next cycle. We have found three such patterns, enclosed by purple squares in Fig. 7: (i) ‘breathing in’ (panel (b)), where three Se2s move close to a central Nb; (ii) ‘breathing out’ (panel (c)), where the three Se2s move close to an empty site; and (iii) Kekule-distortion of a hexagon, where panel (a) illustrates the three phases of a cycle. It has been suggested that Pauling’s resonating valence bond (RVB) of benzene could be a form of dJT effect,[Pauling]. Here we have a circular phonon form of this effect.
@@
Figure 7 illustrates displacement patterns for several $K$-point phonons. Panel (a) shows the Kekule-like pattern in a hexagon of alternating Nb and Se2 pairs (violet box) in the three steps involved in one cycle of mode 6.111This is not strictly a Kekule pattern, as a cycle contains three phases, with alternating short bonds 12-34-56, 23-45-61, and an undistorted hexagon. To confirm our identification of this pattern, we show in Appendix E that a similar pattern arises in graphene, which is a lattice of benzene rings.. Panel (b) shows the breathing-in mode in mode 7, while panel (c) depicts the breathing-out mode of mode 2. The Kekule state is involved in the softening between modes 6 and 7, but it is not the ground state.
The true soft mode lies near 2/3 of $\Gamma-M$. At $M$, all modes are linearly polarized, but as one goes toward $\Gamma$, one set of atoms moves in the orthogonal direction to form an elliptical orbit. Notably, in 1H-NbSe2 an earlier work found two static CDW ground states, corresponding to the breathing-in and breathing-out patterns, with an energy difference of $\sim$8 meV.[NbSe2]
@@
Figure 8: (color online) 2($\Gamma$-M)/3-point phonon displacements. (a) to (c) show modes 8, 6 and 3 respectively in the Y-Z plane, where, ${\bf Z\parallel a_{3}},$ ${\bf X\parallel a_{1}\parallel\Gamma-M}$, and ${\bf Y\perp X,Z}.$ Blue (orange) dots refers to Nb (Se) atoms.(d)Phonon dispersions. Red dots refer to the phonons considered.
@@
VIII.3 Beyond the Born-Oppenheimer (BO) approximation
The EPW code, see Appendix B, includes a perturbative nonadiabatic (i.e., beyond BO) correction, which we have found to be small in NbSe2, although nonadiabatic effects can be much stronger.[nonadi, BBO1, BBO2]
@@ VIII.3 Beyond the Born-Oppenheimer (BO) approximation
Since the BO approximation relies on the large difference in the time scales involved in the electronic and ionic motions, nonadiabatic corrections would be larger when electrons move slowly as in heavy Fermion materials, flat bands, and near van Hove singularities (VHSs). Optical phonons can have strong nonadiabatic effects, as electron hoppings can be strongly enhanced for certain phonon motions, as discussed in the preceding subsection.
VIII.4 Estimating temperature
@@
The broadening parameter $\sigma$ is often introduced as a fictitious temperature to improve the convergence of the DFT self-consistency cycles towards the ground state[fict]. Using the Fermi function with $\sigma=k_{B}T_{\sigma}$, however, leads to a large volume of $k$-space that must be integrated over, so that a more confined broadening function, such as a Gaussian, is a better choice. To gaing a handle on relating $\sigma$ to an effective temperature, we can compare the results of $\sigma$-broadening with direct calculations of the $T$-dependent Lindhard susceptibility, Fig. 9. Focusing on the broadening of the susceptibility peak along $\Gamma-M-K$, it is seen to be very small up to $\sigma=0.016$ Ry ($T_{\sigma}=2526$ K) where the peak is less reduced than for $T=250K$. But, the $\sigma=0.02$ susceptibility ($T_{\sigma}=3157$K) is much more intense and broader than at other $\sigma$ values, with a peak-width that is comparable to the $T=1000$K susceptibility. These results suggest that $T_{\sigma}/T$ is $\leq 10$ and $\geq 3.5$. Since the CDW transition occurs at $\sigma=0.0057$ Ry, or $T_{\sigma}=900$K, we thus estimate that the corresponding $T_{CDW}$ lies between 225K and 90K, which is consistent with the experimental value for monolayer NbSe2, $T_{CDW}=145$K.
@@
Figure 9: (color online) Left panel: Susceptibility using the minimum $\sigma$ value at different Fermi-Dirac temperatures. Right panel: Susceptibility for fixed Fermi-Dirac temperature $T=1$ K for various $\sigma$ values.
IX Conclusions
We show that by accounting for effects of the divergent factor of $\sim 1/\Omega_{q}$ in the self-energy, new insight can be obtained into the nature of the phonon spectra of materials.

### v2

@@ 1.6 A Model of a Kohn ladder
Figure 6 presents illustrative results, which are based on taking the bare phonons from the one-band model solutions for $\sigma=0.00565$ Ry (blue lines in frames (a) and (d)). The results, which illustrate the anticrossing phenomena, are in general accord with the QE results (green line in (f)). Figures 6(a-c) use the clean-limit $\Pi^{\prime}_{33}$, showing that the anomalous nesting peak spoils the agreement with experiment, as expected, while the dirty limit results, Figs. 6(d-f), restore the good agreement. Figs. 6(d-f) also show that for $\alpha\leq 5$, most of the softening is in the LO modes 6 and 8, while the LA mode barely changes. For larger $\alpha$, most of the softening is in the LA mode. The latter regime resembles the QE data for $\sigma=0.03$ Ry (green line in (f)).
@@
Figure 6: (color online) Model calculations of the evolution of dressed frequencies at a series of temperatures scaled by $\alpha$. (a-c) and (d-f) refer to the clean and dirty limits, respectively. For $\alpha=0$ (highest $T$), the one-band solutions for $\omega_{q}$ were calculated using $\sigma=0.03$ Ry (blue lines in (a), (d)). Lower temperatures use full $\Pi^{\prime}_{m}$ matrix. Red curves correspond to $\alpha$ =1 (a), 5 (b), 9 (c), 2.5 (d), 9 (e), and 12 (f). Other curves: green line in (f) = band 3 from $\sigma=0.07$ Ry data; black lines in (c),(f) = band 8 for same $\alpha$ as red lines for $\alpha=9$ without off-diagonal elements, $\Pi^{\prime}_{m,ij}=0$.
@@
Notably, if we turn off mode-coupling, the softening of mode 8 at largest $\alpha$ (black curves in (c),(f)) is too weak for a CDW instability. This can be understood as follows. Considering two modes only, anti-crossing results in the renormalized frequencies given by:
@@
where the $\Omega_{1,2}$ are the renormalized frequencies found from Eqs.  1 (without mode mixing) and $\Omega_{\pm}$ (including mode mixing. Clearly, $\Omega^{2}_{-}$ can go soft completely at the critical mode mixing threshold $\Pi_{12}^{(c)}$ where
@@
which takes place before $\Omega_{1}^{2}$ goes soft (without mode-mode coupling). Off-diagonal self-energies thus generally enhance $T_{CDW}$. Mode-mode mixing can also induce a shift in $\boldsymbol{Q}_{CDW}$. These effects can be shown for NbSe2 by investigating the difference between $\Omega_{\pm}$ and $\Omega_{1,2}$. Note that while the diagonal self-energies of Eq. (1) are real, the off-diagonal terms can be complex. In the irreducible representation of dimension 3, the Hamiltonian determinant contains the term $16\omega_{1}\omega_{2}\omega_{3}Re(\Pi_{12}\Pi_{23}\Pi_{31})$. That is, the phonon energies depend on the relative phases between the off-diagonal components of the self-energy matrix. In contrast, for irreducible representations of lower dimension, the off-diagonal term only enters as $|\Pi_{12}|^{2}$.
@@
Interestingly, the initial electron-phonon interaction at high temperatures is longitudinal, so that the main phonon softening predominantly affects longitudinal modes. However, due to anticrossing, the phonon that softens will appear to be from a transverse mode, unless, as in NbSe2, the softening is along a high-symmetry line or at a point where longitudinal and transverse modes are strictly orthogonal.
1.7 Discussion
1.7.1 Giant Kohn anomalies
We expect strong electron-phonon coupling to often involve optical phonons which strongly modulate electron hopping integrals to yield Kohn anomalies. The resulting Kohn ladder is captured by DFPT, including the resulting CDW whose displacement pattern matches that of the involved optical phonon. (In simulating higher temperatures via $\sigma$, the largest $\sigma$ studied was not large enough to properly restore the large electron-phonon coupling to the highest LO branch, thereby obscuring the origin of the Kohn anomaly.)
@@ 1.7.1 Giant Kohn anomalies
The approach we discussed in this paper should find wide applicability for understanding anticrossing effects and the resulting giant Kohn-ladder anomalies in correlated $d$- and $f$-electron materials. Not including effects of anticrossing could underestimate the strength of electron-phonon coupling. Even in graphene, DFT calculations fail to properly capture the Kohn anomaly52. Notably, a study of spin-lattice relaxation argues that most phonons play no role in relaxation, which is dominated by phonons with rotational or intramolecular vibrational patterns.59 Moreover, even the latter phonons lie at high energies, and only the small tail of these phonons at low energies dominates the relaxation process. Our study suggests that phonons involved in relaxation are all optical, and that the low energy tail is produced by strong electron-phonon interaction in the form of Kohn anomalies, which push some modes to lower energies (soft modes). A chiral Kohn anomaly has been reported in a Weyl semimetal.69
An alternative approach that is not based on using off-diagonal self-energies for extracting bare phonon frequencies72 has been employed to study phonon softening in 1H-TaS27. Notably, Fig. 2(b) of Ref. 7 shows clear evidence of a Kohn ladder remarkably similar to our findings, although the authors state: “The bare phonon dispersion of 1H-TaS2 …, which excludes screening intrinsic to the active subspace, … does not show any Kohn anomalies”, and they thus rule out any role of phonons from other bands. But, the analysis of Ref. 7 overlooks the possibility that phonons in the LA band at low $T$ started out as LO phonons at high $T$ and passed downward via anticrossing, which is the essence of the Kohn ladder effect.
Finally, we note that it was recently reported that the off-diagonal phonon self-energy can also induce anomalous behavior in the electronic dispersion, specifically, a finite-temperature topological-to-trivial transition.42 Then, at the critical temperature, an allowed band crossing appears in the electronic dispersion, similar to the ones we discussed above in the phonon spectra.
@@
Figure 7: (color online) K-point phonon displacements. (a) Mode 6, showing three phases of displacement in one cycle; (b) mode 7, and (c) mode 2 showing a single phase; and (d) phonon dispersions with red dots showing $k$-points used (from top to bottom): mode 6 (a), 7 (b), and 2(c). Yellow boxes highlight one triangle, while violet boxes highlight one hexagon.
1.7.2 Visualizing soft phonons
@@
By using the phonon visualizer software 51, we find that the phonon modes in NbSe2 with significant electron-phonon couplings undergo displacements that are circularly or elliptically polarized, raising the interesting question as to what happens when a circular phonon softens. For a linearly polarized phonon, the frozen phonon CDW will have the maximum distortion associated with the soft phonon. But circular phonons have both amplitude and phase as variables, and the phase variable could lock in uniformly, adopt spatially random fixed directions (as in order-disorder transitions12), or remain free as in a time crystal90, 79. There is also a striking similarity between the circular phonon displacements and the dynamic Jahn-Teller (dJT) phenomena 8, which have been suggested as a source for time crystals. Importantly, many correlated materials are characterized by poorly understood pseudogaps, where “intertwined orders” prevent the establishment of any long-range order. In such a case, a phonon would soften only to a low, but finite frequency, leaving a short-range order with strong fluctuations. While this does not appear to be the case in NbSe2, there are many materials of interest, including cuprates and nickelates, where pseudogap phenomena are observed. If the soft phonons are circular or chiral, the resulting fluctuations would bear a strong resemblance to time crystals, without violating any of the no-go theorems for a time-crystal ground state. It is interesting to consider two $q$-values here: $\bf q=\overrightarrow{\Gamma K}$, which lies in the region where the longitudinal-transverse anticrossing occurs and $\bf q=2/3\overrightarrow{\Gamma M}$, where the soft phonon lies. These two $q$ points are special in that they are 1/3 of a combination of reciprocal lattice vectors. Remarkably, most of the $K$-point phonons are associated with recognizable dJT displacements. One of the earliest dJT systems involves triangular molecules, such as Na3, which can have a Berry phase that leads to an extra orbital angular momentum with half-integer quantization (a ‘time molecule’).57, 20 Similar dJT phenomena can be found on impurity sites in other crystals.
@@
A layer of NbSe2 can be viewed as a lattice of centered triangular molecules, either triangles of Se2 molecules with a central Nb or of Nb molecules with a central Se2. For several $K$-point circular phonons, one cycle of rotation amounts to the central atom displacing sequentially toward each of the triangular atoms, leading to a time-averaged symmetrical pattern (yellow squares in Fig. 7), as in Refs. 57, 20. However, a lattice of dJT molecules has extra degrees of freedom, which also have a dJT interpretation. On each of the three steps in a cycle, a pattern forms on one set of atoms, then dissolves and reforms on a second and then a third set before it repeats in the next cycle. We have found three such patterns, enclosed by purple squares in Fig. 7: (i) ‘breathing in’ (panel (b)), where three Se2s move close to a central Nb; (ii) ‘breathing out’ (panel (c)), where the three Se2s move close to an empty site; and (iii) Kekulé-distortion of a hexagon, where panel (a) illustrates the three phases of a cycle.
In connection with circular phonons, it is interesting to recall that one would intuitively expect a benzene ring (C6 molecule) to have alternating single and double bonds, where the two carbons joined by a double bond are slightly closer to each other. However, experimentally all the carbons in the benzene ring seem to be equivalent. Kekulé interpreted this via a dynamic model of benzene, where the long and short bonds interchange dynamically in cycles of oscillations.73 This idea was formalized in Pauling’s resonating valence bond (RVB) of benzene, which can be considered a form of dJT effect73. In our case, we have a circular phonon version of this effect.
@@
Figure 7 illustrates displacement patterns for several $K$-point phonons. Panel (a) shows the Kekulé-like pattern in a hexagon of alternating Nb and Se2 pairs (violet box) in the three steps involved in one cycle of mode 6. (This is not strictly a Kekulé pattern, as a cycle contains three phases, with alternating short bonds 12-34-56, 23-45-61, and an undistorted hexagon. To confirm our identification of this pattern, we show in SI-2 that a similar pattern arises in graphene, which is a lattice of benzene rings.) Panel (b) shows the breathing-in mode in mode 7, while panel (c) depicts the breathing-out mode of mode 2. The Kekulé state is involved in the softening between modes 6 and 7, but it is not the ground state. The true soft mode lies near 2/3 of $\Gamma-M$. At $M$, all modes are linearly polarized, but as one goes toward $\Gamma$, one set of atoms moves in the orthogonal direction to form an elliptical orbit. Notably, in 1H-NbSe2, an earlier work found two static CDW ground states, corresponding to the breathing-in and breathing-out patterns, with an energy difference of $\sim$8 meV.14
@@
Figure 8: (color online) 2($\Gamma$-M)/3-point phonon displacements. (a) to (c) show modes 8, 6 and 3, respectively, in the Y-Z plane, where $\bf Z\parallel a_{3}$}, $\bf X\parallel a_{1}\parallel\Gamma-M$}, and $\bf Y\perp X,Z$}. Blue (orange) dots refer to Nb (Se) atoms. (d)Phonon dispersions. Red dots refer to the phonons considered.
@@
1.7.3 Beyond the Born-Oppenheimer (BO) approximation
@@
The EPW code, see LABEL:SM3, includes a perturbative nonadiabatic (i.e., beyond BO) correction, which we have found to be small in NbSe2, although nonadiabatic effects can be much stronger.58, 5, 74 Since the BO approximation relies on the large difference in the time scales involved in the electronic and ionic motions, nonadiabatic corrections would be larger when electrons move slowly as in heavy Fermion materials, flat bands, and near van Hove singularities (VHSs). Optical phonons can have strong nonadiabatic effects, as electron hoppings can be strongly enhanced for certain phonon motions, as discussed in the preceding subsection.
1.7.4 Estimating temperature
@@
The broadening parameter $\sigma$ is often introduced as a fictitious temperature to improve the convergence of the DFT self-consistency cycles towards the ground state22. Using the Fermi function with $\sigma=k_{B}T_{\sigma}$, however, leads to a large volume of $k$-space that must be integrated over, so that a more confined broadening function, such as a Gaussian, is a better choice. To gain a handle on relating $\sigma$ to an effective temperature, we can compare the results of $\sigma$-broadening with direct calculations of the $T$-dependent Lindhard susceptibility, Fig. 9. Focusing on the broadening of the susceptibility peak along $\Gamma-M-K$, it is seen to be very small up to $\sigma=0.016$ Ry ($T_{\sigma}=2526$ K) where the peak is less reduced than for $T=250K$. But, the $\sigma=0.02$ susceptibility ($T_{\sigma}=3157$K) is much more intense and broader than at other $\sigma$ values, with a peak-width that is comparable to the T=1000T=1000K susceptibility. These results suggest that $T_{\sigma}/T$ is $\leq 10$ and $\geq 3.5$. Since the CDW transition occurs at $\sigma=0.0057$ Ry, or $T_{\sigma}=900$K, we thus estimate that the corresponding $T_{CDW}$ lies between 225K and 90K, which is consistent with the experimental value for monolayer NbSe2, $T_{CDW}=145$K.
@@
Figure 9: (color online) Left panel: Susceptibility using the minimum $\sigma$ value at various Fermi-Dirac temperatures. Right panel: Susceptibility for fixed Fermi-Dirac temperature $T=1$ K for various $\sigma$ values.
2 Conclusions